HUBBLE SPACE TELESCOPE DAILY REPORT #2112 PERIOD COVERED: 0000Z (UTC) 04/28/98 - 0000Z (UTC) 04/29/98 Daily Status Report as of 119/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Three Sets of WF/PC-2 7712 (WF/PC-2 Cycle 7 Supplemental Darks pt2/3) The WF/PC-2 was used to make a series of Cycle 7 supplemental calibration darks. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed STIS/CCD 7926 (CCD Dark and Bias Monitor -- Continued) The Space Telescope Imaging Spectrograph (CCD) was used to make several dark and bias calibration observations. This is to measure CCD dark current and bias rates for the two supported amplifier settings {gain = 1 and gain = 4} in order to gather data for making superdarks and superbiases. These observations will also track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed NIC/2 7331 (Near-IR Properties of the Bulges of Spiral Galaxies) The Near Infrared Camera (Camera Two) was used to make infrared observations of the bulge in spiral galaxy NGC-3169. Near-IR data at HST resolution are essential for correcting for the effect of dust and will enable us to derive robust nuclear properties for the sample. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed NIC/2 7330 (The Fueling of Active Nucleii: A NICMOS Snapshot Survey) The Near Infrared Camera (Camera Two) was used to probe the dark matter distribution of NGC-4260 far from the center by observing intervening absorption from the MgII doublet (lambda 2800) in the spectrum of one background QSO located in a geometrically optimal position. HST observations are required because this feature is in a wavelength range not accessible from the ground, and the QSO magnitudes and expected strength of absorption make IUE observations unfeasible. The proposal completed as expected with no anomalies. 1.5 Completed STIS/CCD 7803 (Daily Darks to Update Acquisition Bad Pixel Table Part III) The Space Telescope Imaging Spectrograph (CCD) was used to make daily dark calibration updates. These daily darks are intended to identify hot pixels for update to the acquisition on-board FSW hot pixel table. Daily hot pixel lists can also be used for science data analysis. The observations were completed as planned, and no anomalies were reported. 1.6 Completed Two Sets of NIC/1 7232 (NICMOS Spectrophotometry and High Resolution Imaging of HD 98800) The Near Infrared Camera (Camera One) was used to make spectrophotometry and high resolution imaging of star HD 98800. This star system is located nearby at a probable distance of 20 pcs. It is a unique stellar system consisting of two K7-V stars separated by 0.8 arcsec. A dark calibration image was also taken. The observations were executed as planned, and no problems were encountered. 1.7 Completed STIS/MA1 7401 (A Critical Test of Crystallization in White Dwarf Stars) The Space Telescope Imaging Spectrograph (MA1) was used to make observations of the white dwarf star BPM37093. Theoretical calculations indicate that the massive pulsating DA white dwarf, BPM37093 has a crystallized ion interior. A wavelength calibration image was also taken. The observations were executed as scheduled, and no anomalies were noted. 1.8 Completed Five Sets of WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. The observations were completed as planned, and no problems were reported. 1.9 Completed Four Sets of NIC/2/3 7907 (NICMOS Pure Parallel Winter 97-98) The Near Infrared Camera (Cameras Two and Three) was used to make pure parallel spectrographic observations of sky regions when another instrument is prime. The observations were completed as planned, and no anomalies were reported. 1.10 Completed WF/PC-2/STIS/CCD/NIC/1/2 7307 (The Formation and Evolution of Rich Star Clusters in the LMC) The WF/PC-2, the Space Telescope Imaging Spectrograph (CCD), and the Near Infrared Camera (Cameras One and Two) were used to make observations of the rich star cluster NGC-1805 which is located within the Large Magellanic Cloud in order to study its formation and evolution. These observations will help us understand the origin and evolution of rich star clusters in our own galaxy. The observations were completed as scheduled, and no anomalies were reported. 1.11 Completed STIS/MA2 7604 (Cycle 7 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA2) was used to make MAMA dark measurements. These darks will be used in later image reductions and also trended over the instrument's life time as an indicator of the detector's degradation or damage. The observations were completed as scheduled, and no problems were reported. 1.12 Completed WF/PC-2 7277 (Calibration of the Cepheid P-L Relation with Observations of the Maser-Host Galaxy NGC) The WF/PC-2 was used to make observations of the maser-host galaxy NGC-4258 to study its Cepheid variable stars. These data will help refine the calibration of the Cepheid period-luminosity relationship. The observations were completed as scheduled, and no problems were reported. 1.13 Completed Two Sets of STIS/CCD 7908 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal) The Space Telescope Imaging Spectrograph (CCD) was used to make calibration observations as part of the POMS Test Proposal. This was a STIS non-scripted parallel proposal. The observations were completed as planned, and no problems were reported. 1.14 Completed WF/PC-2 7590 (Cosmological Parameters Omega and Lambda From High-Redshift Type IA Supernovae) The WF/PC-2 was used to make observations of type 1A supernova 14H0.9Z1 in order to study its redshift. The HST was used to make direct measurements of the cosmological parameters Omega, Lambda, and curvature using Type 1A supernovae as a calibrated "standard candle". The observations were executed as scheduled, and no anomalies were noted. 1.15 Completed NIC/1 7867 (NICMOS Imaging of the Dusty CfA Seyfert Nuclei) The Near Infrared Camera (Camera One) was used to make observations of the Seyfert Galaxy NGC-5033. These data will allow us to determine if the Seyfert Type 1/2 dichotomy is best explained by the unified model, the evolutionary picture, or a combination of both of these scenarios. As described in 2.1, the acquisition for this proposal defaulted to fine lock backup on one FGS only. Otherwise, the observations were executed as scheduled, and no further anomalies were reported. 1.16 Completed NIC/1/2 7849 (Metallicity and the Cepheid Distance Scale) The Near Infrared Camera (Cameras One and Two) was used to make observations of Cepheid variable stars in the M101-IB galaxy. The local distance scale to galaxies is largely based on the Cepheid period-luminosity relation. Motivated by the goal of measuring an accurate value of the Hubble constant, HST has been used successfully to discover scores of Cepheid variables in each of about 20 galaxies. The observations were executed as scheduled, and no anomalies were noted. 1.17 Completed Two sets of STIS/CCD 7912 (STIS Parallel Archive Proposal - Nearby Galaxies - Imaging and Spectroscopy) The Space Telescope Imaging Spectrograph (CCD) was used to make parallel observations of nearby galaxies. This survey will be useful to study the star formation histories, chemical evolution, and distances to these galaxies. These data will be placed immediately into the Hubble Data Archive. The observations were completed as scheduled, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 7 Successful : 7 The acquisition at 118/132247Z defaulted to fine lock backup on FGS-3 only. HSTAR 6452 was written. The proposal described in 1.15 may have been affected. Scheduled Reacquisitions : 8 Successful : 8 2.2 FHST Updates: Scheduled : 13 Successful : 13 2.3 Operations Notes: An RMGA calibration was performed at 118/2257Z. Using ROP SR-1, the SSR correctable EDAC counts were cleared three times. Flight operations personnel successfully transferred to the SAMS backup facility at 119/0700Z, and will remain there throughout the day. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH